# TractShapeNet: Efficient Multi-Shape Learning with 3D Tractography Point Clouds

### v1

@@ Abstract
Brain imaging studies have demonstrated that diffusion MRI tractography geometric shape descriptors can inform the study of the brain’s white matter pathways and their relationship to brain function. In this work, we investigate the possibility of utilizing a deep learning model to compute shape measures of the brain’s white matter connections. We introduce a novel framework, TractShapeNet, that leverages a point cloud representation of tractography to compute five shape measures: length, span, volume, total surface area, and irregularity. We assess the performance of the method on a large dataset including 1065 healthy young adults. Experiments for shape measure computation demonstrate that our proposed TractShapeNet outperforms other point-cloud-based neural network models in both the Pearson correlation coefficient and normalized error metrics. We compare the inference runtime results with the conventional shape computation tool DSI-Studio. Our results demonstrate that a deep learning approach enables faster and more efficient shape-measure computation. We also conduct experiments on two downstream language cognition prediction tasks, showing that shape measures from TractShapeNet perform similarly to those computed by DSI-Studio. Our code will be available at: https://github.com/SlicerDMRI/TractSh-apeNet.
@@ 1 Introduction
In recent years, deep learning methods have been successful in dMRI tractography analysis [12], but computing shape measures in a way that can best take advantage of deep networks remains an open challenge. Deep learning using point clouds is promising for the analysis of geometric structures in fields like 3D object retrieval and medical imaging [10]. A point cloud is a finite set of points with 3D coordinates and optional attributes such as normals or features. In dMRI tractography, white matter connections are represented as streamlines that consist of sequences of points (Fig. 1). Recent works have demonstrated the potential of point-based neural networks for the analysis of dMRI tractography in tasks such as parcellation, filtering, and prediction of cognitive performance [28, 7, 2]
@@ 2 Methods
This study evaluates our proposed method on the large-scale Human Connectome Project Young Adult (HCP-YA) dataset, which includes 1065 healthy young adults (575 females and 490 males, 28.7 years old on average) [26, 25]. Whole brain tractography is generated from each subject’s dMRI data using a two-tensor unscented Kalman filter method [18]. Tractography is then parcellated into individual fiber clusters using an anatomically curated tractography brain atlas [30]. We focus on the key task of shape computation for association pathways [29]. We employ fiber clusters within ten left hemisphere association tracts of arcuate fasciculus (AF), cingulum bundle (CB), extreme capsule (EmC), inferior longitudinal fasciculus (ILF), middle longitudinal fasciculus (MdLF), and uncinate fasciculus (UF). This results in 73 fiber clusters per subject, for a total of 77,745 clusters across the 1,065 subjects for training and testing our methods.
@@ 2.2 Shape Measures
We compute ground truth shape measures for each fiber cluster using DSI-Studio [29]. Detailed definitions of the shape measures can be seen in [29, 16]. We select five shape descriptors for analysis in this study: length, span, volume, total surface area, and irregularity. Length is the average streamline length of the fiber cluster, span is the distance between the two ends of the fiber cluster, and volume is the volume of voxels occupied by the fiber cluster. We include total surface area and irregularity because we have shown that they are highly informative for the prediction of individual cognitive performance [16]. Total surface area is the area of the outermost or surface voxels occupied by the fiber cluster, while irregularity quantifies how different the fiber cluster shape is from an idealized cylinder.
@@ 2.3 Representing Fiber Clusters as Point Clouds
Our model is trained using the right-anterior-superior (RAS) coordinates coordinate points of individual fiber clusters as independent (input) variables, with the shape measures as the dependent (output) variables. Following TractGeoNet [7], we represent each fiber cluster as a point cloud. We construct 73 point clouds for every subject, corresponding to the 73 fiber clusters. We leverage a random sampling strategy to select N random points from each cluster as input for the deep learning model. Random sampling has previously been shown to be an effective strategy for the analysis of tractography as point clouds [7] because this strategy achieves an equal probability of selection across all points [21].
@@ 2.4 Network Architecture and Proposed Loss Function
We introduce a hybrid loss function $L_{total}$ that combines the subnetwork loss functions ($L_{1}$, $L_{2}$) with a proposed weighted pairwise Siamese-Fourier loss function ($L_{SF}$) to optimize the model performance as shown in Equation (1). The subnetwork loss functions calculate the error between the output and the ground truth. The weighted pairwise loss function $L_{SF}$ complements this by capturing patterns across a Fourier frequency domain. As shown in Equation (2), we utilize the discrete Fourier transformation (DFT) to transform our predicted output and ground truth shape measures into Fourier frequency domain representations [8, 19]. This transform can potentially enhance multi-shape learning by leveraging global relationships across the different shape measures. $e^{-\tfrac{j2\pi}{N}kn}$ is Euler’s formula that transforms the Fourier frequency domain, where $k$ and $n$ represent the $k$th frequency component and the $n$th sample, respectively. $L_{SF}$ calculates the mean squared error between the output differences ($O_{1}-O_{2}$) and corresponding ground truth differences ($GT_{1}-GT_{2}$) in Fourier frequency space:
@@ 2.5 Implementation Details
The model is trained and evaluated with a train test split of 80% and 20%. The model is set with a batch size of 128 for 200 epochs. The initial learning rate is 0.0005 with the Adam optimizer [13] with a weight decay of 0.005. The scheduler updates the learning rate with a decay factor (gamma) of 0.1 every 200 steps. We tuned the weight ($\alpha$) of the Siamese Fourier loss function loss and set it to 3 for optimal performance. The code is implemented using PyTorch v1.13 [20]. All experiments are conducted on the Jetstream2 cloud computing environment, configured with an NVIDIA A100 40 GB GPU, 117 GB of RAM, and 32 CPU cores [11, 3].
@@ 3.1 Evaluation Metrics
We employ the Pearson correlation coefficient (Pearson’s $r$) [23] and the normalized mean squared error (nMSE) to evaluate model performance. Pearson’s r measures the strength and direction (positive or negative) of the correlation between predicted and ground-truth shape measurements. nMSE measures the error between the predicted and ground truth values [15, 14]. nMSE is normalized to a range between 0 and 1, where an nMSE value close to 0 represents a low error when compared to the ground truth.
@@ 3.3 Evaluation of Inference Runtime
We calculate the average inference runtime in milliseconds (ms) to compute the five shape measures for individual fiber clusters. Results in Table 3 show that deep point cloud models show faster inference runtimes than DSI-Studio. Although TractShapeNet is slightly less efficient than TractGeoNet, it offers a tradeoff with improved performance across all shape measures.
@@
Methods DSI-Studio [29] PointNet [5] TractGeoNet [7] TractShapeNet
Runtime 147.97 ± 77.0355 ms 1.44 ± 0.0003 ms 0.89 ± 0.0001ms 1.24 ± 0.0004ms

### v4

@@ Abstract
Brain imaging studies have demonstrated that diffusion MRI tractography geometric shape descriptors can inform the study of the brain’s white matter pathways and their relationship to brain function. In this work, we investigate the possibility of utilizing a deep learning model to compute shape measures of the brain’s white matter connections. We introduce a novel framework, TractShapeNet, that leverages a point cloud representation of tractography to compute five shape measures: length, span, volume, total surface area, and irregularity. We assess the performance of the method on a large dataset including 1,065 healthy young adults. Experiments for shape measure computation demonstrate that our proposed TractShapeNet outperforms other point-cloud-based neural network models in both the Pearson correlation coefficient and normalized error metrics. We compare the inference runtime results with the conventional shape computation tool DSI-Studio. Our results demonstrate that a deep learning approach enables faster and more efficient shape-measure computation. We also conduct experiments on two downstream language cognition prediction tasks, showing that shape measures from TractShapeNet perform similarly to those computed by DSI-Studio. Our code will be available at: https://github.com/SlicerDMRI/TractSh-apeNet.
@@ 1 Introduction
In recent years, deep learning methods have been successful in dMRI tractography analysis [12], but computing shape measures in a way that can best take advantage of deep networks remains an open challenge. Deep learning using point clouds is promising for the analysis of geometric structures in fields like 3D object retrieval and medical imaging [10]. A point cloud is a finite set of points with 3D coordinates and optional attributes such as normals or features. In dMRI tractography, white matter connections are represented as streamlines that consist of sequences of points (Fig. 1). Recent works have demonstrated the potential of point-based neural networks for the analysis of dMRI tractography in tasks such as parcellation, filtering, and prediction of cognitive performance [28, 7, 2].
@@ 2 Methods
This study evaluates our proposed method on the large-scale Human Connectome Project Young Adult (HCP-YA) dataset, which includes 1,065 healthy young adults (575 females and 490 males, 28.7 years old on average) [26, 25]. Whole brain tractography is generated from each subject’s dMRI data using a two-tensor unscented Kalman filter method [18]. Tractography is then parcellated into individual fiber clusters using an anatomically curated tractography brain atlas [30]. We focus on the key task of shape computation for association pathways [29]. We employ fiber clusters within ten left hemisphere association tracts of arcuate fasciculus (AF), cingulum bundle (CB), extreme capsule (EmC), inferior longitudinal fasciculus (ILF), middle longitudinal fasciculus (MdLF), and uncinate fasciculus (UF). This results in 73 fiber clusters per subject, for a total of 77,745 clusters across the 1,065 subjects for training and testing our methods.
@@ 2.2 Shape Measures
We compute ground-truth shape measures for each fiber cluster using DSI-Studio [29]. Detailed definitions of the shape measures can be seen in [29, 16]. We select five shape descriptors for analysis in this study: length, span, volume, total surface area, and irregularity. Length is the average streamline length of the fiber cluster, span is the distance between the two ends of the fiber cluster, and volume is the volume of voxels occupied by the fiber cluster. We include total surface area and irregularity because we have shown that they are highly informative for the prediction of individual cognitive performance [16]. Total surface area is the area of the outermost or surface voxels occupied by the fiber cluster, while irregularity quantifies how different the fiber cluster shape is from an idealized cylinder.
@@ 2.3 Representing Fiber Clusters as Point Clouds
Our model is trained using the right-anterior-superior (RAS) coordinate points of individual fiber clusters as independent (input) variables, with the shape measures as the dependent (output) variables. Following TractGeoNet [7], we represent each fiber cluster as a point cloud. We construct 73 point clouds for every subject, corresponding to the 73 fiber clusters. We leverage a random sampling strategy to select $N$ random points from each cluster as input for the deep learning model. Random sampling has previously been shown to be an effective strategy for the analysis of tractography as point clouds [7] because this strategy achieves an equal probability of selection across all points [21].
@@ 2.4 Network Architecture and Proposed Loss Function
We introduce a hybrid loss function $L_{total}$ that combines the subnetwork loss functions ($L_{1}$, $L_{2}$) with a proposed weighted pairwise Siamese-Fourier loss function ($L_{SF}$) to optimize the model performance as shown in Equation (1). The subnetwork loss functions calculate the error between the output and the ground-truth. The weighted pairwise loss function $L_{SF}$ complements this by capturing patterns across a Fourier frequency domain. As shown in Equation (2), we utilize the discrete Fourier transformation (DFT) to transform our predicted output and ground-truth shape measures into Fourier frequency domain representations [8, 19]. This transform can potentially enhance multi-shape learning by leveraging global relationships across the different shape measures. $e^{-\tfrac{j2\pi}{N}kn}$ is Euler’s formula that transforms the Fourier frequency domain, where $k$ and $n$ represent the $k$th frequency component and the $n$th sample, respectively. $L_{SF}$ calculates the mean squared error between the output differences ($O_{1}-O_{2}$) and corresponding ground-truth differences ($GT_{1}-GT_{2}$) in Fourier frequency space:
@@ 2.5 Implementation Details
The model is trained and evaluated with a train test split of 80% and 20%. The model is set with a batch size of 128 for 200 epochs. The initial learning rate is 0.0005 with the Adam optimizer [13] with a weight decay of 0.005. The scheduler updates the learning rate with a decay factor (gamma) of 0.1 every 200 steps. We tuned the weight ($\alpha$) of the Siamese Fourier loss function and set it to 3 for optimal performance. The code is implemented using PyTorch v1.13 [20]. All experiments are conducted on the Jetstream2 cloud computing environment, configured with an NVIDIA A100 40 GB GPU, 117 GB of RAM, and 32 CPU cores [11, 3].
@@ 3.1 Evaluation Metrics
We employ the Pearson correlation coefficient (Pearson’s $r$) [23] and the normalized mean squared error (nMSE) to evaluate model performance. Pearson’s $r$ measures the strength and direction (positive or negative) of the correlation between predicted and ground-truth shape measurements. nMSE measures the error between the predicted and ground-truth values [15, 14]. nMSE is normalized to a range between 0 and 1, where an nMSE value close to 0 represents a low error when compared to the ground-truth.
@@ 3.3 Evaluation of Inference Runtime
We calculate the average inference runtime in milliseconds (ms) to compute the five shape measures for individual fiber clusters. Results in Table 3 demonstrate that deep point cloud models show faster inference runtimes than DSI-Studio. Although TractShapeNet is slightly less efficient than TractGeoNet, it offers a tradeoff with improved performance across all shape measures.
@@
Methods DSI-Studio [29] PointNet [5] TractGeoNet [7] TractShape-Net
Runtime 41.10 ±21.3987 1.44 ± 0.0003 ms 0.89 ± 0.0001ms 1.24 ± 0.0004ms